OpenStudy (anonymous):

1. Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (anonymous):

do u know how to do it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

can i ask what grade you are in just to be sure

OpenStudy (anonymous):

um yea 9

OpenStudy (agent47):

I have no idea what they want from you, but I'm just going to solve that equation: 2.5x – 10.5 = 64(0.5x) 2.5x - 10.5 = 32x 2.5x - 32x = 10.5 -29.5x = 10.5 x = -10.5/29.5

OpenStudy (anonymous):

yea that to solve it the just wanna see the steps in order so thx a lot happy new year!!!!! thx agent 47